OpenStudy (anonymous):

i have 3 more can u help with them?

OpenStudy (anonymous):

@DatChinookGuy & @jackmullen55

OpenStudy (anonymous):

Yeah I can

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

The angles shown below are supplementary: A pair of supplementary angles is shown. One angle measures 10x and the other angle measures 70 degrees. What is the value of x? 10° 11° 20° 21° Question 10 (Multiple Choice Worth 5 points) (05.02 MC) Based on the figure below, what is the value of x? A right angle is shown divided in two parts. The measure of the angle of one part is 30 degrees and the measure of the other part is 5x plus 15 degrees. 3 9 12 15

OpenStudy (anonymous):

Do you have the picture?

OpenStudy (anonymous):

Nvm don't need it

OpenStudy (anonymous):

So supplementary angles add up to 180. 180 - 70 = 110 So we know that 10x = 110. what is x?

OpenStudy (anonymous):

10? im not sure

OpenStudy (anonymous):

No its 11. because 10 x 11 equals 110

OpenStudy (anonymous):

ok thanks what about the second one

OpenStudy (anonymous):

its 3 isnt it

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Yes I think so

OpenStudy (anonymous):

ok @DatChinookGuy do u think for the second one its 3? im just making sure im right

OpenStudy (anonymous):

@lovesvd10 yeah after further checking and looking around he's right, second one is definitely 3.

OpenStudy (anonymous):

yes thank one there is just one more and im happy u guys can help me bc im terrible at math

OpenStudy (anonymous):

The figure below shows a shaded circular region inside a larger circle: A shaded circle is shown inside another larger circle. The radius of the smaller circle is labeled as r and the radius of the larger circle is labeled as R. On the right side of the image is written r equal to 2 inches and below r equal to 2 inches is written R equal to 5 inches. What is the probability that a point chosen inside the larger circle is not in the shaded region? 84% 50% 42% 16%

OpenStudy (anonymous):

if i dont pass these exams then im doing summer school and i dont want that

OpenStudy (anonymous):

I don't want that either

OpenStudy (anonymous):

Could I get the picture for this?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Noticing the 2 inches below r equal to 2 inches is written equal to 5 inches, wtf? lol. "probability that a point chosen inside the larger circle is not in the shaded region?" ANYWAYS, depending on the equality percentages I'd say D, 16%. Only because the way the image is showing off to the right. The 2inches is 1/2 OP, Hope this helps, have a nice day :) Good luck on your exam (; Tag me on anymore you need help with in the future.

OpenStudy (anonymous):

yes it helps and thanks to both of u and i will and are you good with civics...i need help on that once i get the password

OpenStudy (anonymous):

@DatChinookGuy could u give @jackmullen55 a medal .. u both deserve one!

OpenStudy (anonymous):

Yes im good with both civics, math civics, and Honda Civic SI car mechanics. ;) And sure I will.

OpenStudy (anonymous):

but yes also good with the study of the rights and duties of citizenship. if you need help with that as well.

OpenStudy (anonymous):

ok..i forgot this one im not sure on th answer i think its a but im nto sure

OpenStudy (anonymous):

An architect is creating a scale drawing of a school computer lab. The length of the lab is 36 feet and the width of the lab is 48 feet. If each 12 feet of the lab equals 2 centimeters on a scale drawing, which of the following drawings is the scale drawing of the computer lab? A rectangle is shown. The length of the rectangle is labeled as length equal to 6 cm and the width is labeled as width equal to 8 cm. A rectangle is shown. The length of the rectangle is labeled as length equal to 3 cm and the width is labeled as width equal to 4 cm. A rectangle is shown. The length of the rectangle is labeled as length equal to 12 cm and the width is labeled as width equal to 24 cm. A rectangle is shown. The length of the rectangle is labeled as length equal to 24 cm and the width is labeled as width equal to 36 cm.

OpenStudy (anonymous):

What no medal.... :'(

OpenStudy (anonymous):

Piece of cake, we can simply eliminate the answers, NOT: A rectangle is shown. The length of the rectangle is labeled as length equal to 6 cm and the width is labeled as width equal to 8 cm. NOT: A rectangle is shown. The length of the rectangle is labeled as length equal to 12 cm and the width is labeled as width equal to 24 cm. NOT:A rectangle is shown. The length of the rectangle is labeled as length equal to 24 cm and the width is labeled as width equal to 36 cm. We can obviously tell the answer is: A rectangle is shown. The length of the rectangle is labeled as length equal to 3 cm and the width is labeled as width equal to 4 cm. ( B is your answer )

OpenStudy (anonymous):

@jackmullen55 I gave you a medal so we both got a medal.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Yay! Thanks man @DatChinookGuy

OpenStudy (anonymous):

-_- i choose b the first time wow just wow

OpenStudy (anonymous):

@lovesvd10 haha see your smart (; @jackmullen55 no problem man, nice job.

OpenStudy (anonymous):

:D thanks an both of u did a great job im so happy u guys could help me out! i swear if i could id hug both of u!<3

OpenStudy (anonymous):

Haha thanks. BTW what grade are you in 10th? This is some hi level stuff

OpenStudy (anonymous):

im about to be in 8th

OpenStudy (anonymous):

Oh cool. So your 7th.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You look a bit older in your picture i suppose. I'm in 9th

OpenStudy (anonymous):

im dead!!!! im back down to a D omg and thats not me

OpenStudy (anonymous):

What!?!?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

i got 33.33% on the damn exam omg

OpenStudy (anonymous):

Uh oh... Did you get the ones I answered wrong?

OpenStudy (anonymous):

111 08.13 Segment Two Exam Part One Segment: 2 25 / 75 33.33 %

OpenStudy (anonymous):

Do you go to Laurel Springs School?

OpenStudy (anonymous):

no flvs

OpenStudy (anonymous):

Question 1 (Worth 5 points) (05.01 MC) An architect is creating a scale drawing of a school computer lab. The length of the lab is 36 feet and the width of the lab is 48 feet. If each 12 feet of the lab equals 2 centimeters on a scale drawing, which of the following drawings is the scale drawing of the computer lab? A rectangle is shown. The length of the rectangle is labeled as length equal to 6 cm and the width is labeled as width equal to 8 cm. A rectangle is shown. The length of the rectangle is labeled as length equal to 3 cm and the width is labeled as width equal to 4 cm. A rectangle is shown. The length of the rectangle is labeled as length equal to 12 cm and the width is labeled as width equal to 24 cm. A rectangle is shown. The length of the rectangle is labeled as length equal to 24 cm and the width is labeled as width equal to 36 cm. Points earned on this question: 0 Question 2 (Worth 5 points) (06.04 MC) Stephen has a counter that is orange on one side and brown on the other. The counter is shown below: A circular counter is shown. The top surface of the counter is shaded in a lighter shade of gray and Orange is written across this section. The bottom section of the counter is shaded in darker shade of gray and Brown is written across it. Stephen flips this counter 24 times. What is the probability that the 25th flip will result in the counter landing on orange side up? fraction 24 over 25 fraction 1 over 24 fraction 1 over 4 fraction 1 over 2 Points earned on this question: 5 Question 3 (Worth 5 points) (05.04 MC) A mirror frame in the shape of an oval is shown below. The ends of the frame form semicircles: An oval is formed by a rectangle with semicircles at each end. The length of the rectangle is 62 inches. The width of the rectangle is 27 inches. Which of the following is the perimeter of the inner edge of the frame? (π = 3.14) 146.78 inches 208.78 inches 293.56 inches 347.56 inches Points earned on this question: 5 Question 4 (Worth 5 points) (05.04 LC) The image below is a triangle drawn inside a circle with center O: A triangle is shown inscribed inside a circle. The leg of the triangle labeled 6 inches passes through the center of the circle, O. The other two legs are labeled as 4 inches and 3 inches. Which of the following expressions shows the area, in square inches, of the circle? (π = 3.14) 3 ⋅ 3.14 ⋅ 22 3.14 ⋅ 32 3.14 ⋅ 22 3.14 ⋅ 3 Points earned on this question: 0 Question 5 (Worth 5 points) (06.02 HC) Jamie has a deck of 60 sports cards, of which some are baseball cards and some are football cards. Jamie pulls out a card randomly from the deck, records its type, and replaces it in the deck. Jamie has already recorded six baseball cards and nine football cards. Based on these data, what is, most likely, the number of baseball cards in the deck? 12 15 24 30 Points earned on this question: 0 Question 6 (Worth 5 points) (05.01 LC) The current size of an image on Sandy's computer is shown below: A rectangle is shown. The length of the rectangle is labeled as length equal to 4 inches and the width is labeled as width equal to 6 inches. What are the dimensions of the image at fraction 1 over 2 times its current size? Length = 8 inches, width = 12 inches Length = 6 inches, width = 8 inches Length = 2 inches, width = 4 inches Length = 2 inches, width = 3 inches Points earned on this question: 0 Question 7 (Worth 5 points) (05.05 LC) The base length of a triangle is 8 feet and the height is 4 feet. What is the area of the triangle? 4 square feet 8 square feet 16 square feet 32 square feet Points earned on this question: 0 Question 8 (Worth 5 points) (05.05 MC) The area of a triangle is 24 square inches. What is the height of the triangle if the base length is 8 inches? 6 inches 8 inches 12 inches 16 inches Points earned on this question: 0 Question 9 (Worth 5 points) (05.02 MC) The angles shown below are supplementary: A pair of supplementary angles is shown. One angle measures 10x and the other angle measures 70 degrees. What is the value of x? 10° 11° 20° 21° Points earned on this question: 5 Question 10 (Worth 5 points) (05.02 MC) Based on the figure below, what is the value of x? A right angle is shown divided in two parts. The measure of the angle of one part is 30 degrees and the measure of the other part is 5x plus 15 degrees. 3 9 12 15 Points earned on this question: 0 Question 11 (Worth 5 points) (05.05 LC) A wooden block in the shape of a rectangular pyramid is shown below: A shaded right rectangular pyramid is shown. If a cross section of the block is cut in a plane parallel to the base of the pyramid, which of the following shapes describes the cross section? Triangle Rectangle Trapezoid Hexagon Points earned on this question: 0 Question 12 (Worth 5 points) (05.02 LC) The lengths of two sides of a triangle are 10 inches and 4 inches. Which of the following dimensions is the third side of this triangle? 4 inches 5 inches 6 inches 7 inches Points earned on this question: 0 Question 13 (Worth 5 points) (06.02 LC) A coin is flipped 70 times and lands on tails 21 times. What is the relative frequency of the coin landing on tails? 30% 50% 60% 70% Points earned on this question: 5 Question 14 (Worth 5 points) (06.04 MC) The figure below shows a shaded circular region inside a larger circle: A shaded circle is shown inside another larger circle. The radius of the smaller circle is labeled as r and the radius of the larger circle is labeled as R. On the right side of the image is written r equal to 2 inches and below r equal to 2 inches is written R equal to 5 inches. What is the probability that a point chosen inside the larger circle is not in the shaded region? 84% 50% 42% 16% Points earned on this question: 0 Question 15 (Worth 5 points) (06.01 LC) During the grand opening of an electronics store, any customer has a 17% chance of winning a free laptop. What is the chance that a customer does not win a free laptop? 81% 83% 51% 37% Points earned on this question: 5

OpenStudy (anonymous):

my mom is totally going to make me do summer school ugh

OpenStudy (anonymous):

That sucks! I'm so sorry!

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

68.42 (D)

OpenStudy (anonymous):

my teacher has to grade 4 essays that are worth 40 points so maybe ill be back in a C soon

OpenStudy (anonymous):

Fingers crossed

OpenStudy (anonymous):

yes very!

OpenStudy (anonymous):

That's a lot of essays tho

OpenStudy (anonymous):

ik but it was easy

OpenStudy (anonymous):

what did you have to write about?

OpenStudy (anonymous):

Alfred draws candles randomly from a pack containing 4 colored candles of the same shape and size. There are 2 red candles, 1 green candle, and 1 blue candle. He draws 1 candle and then draws another candle without replacing the first one. Find the probability of picking 1 red candle followed by another red candle, and show the equation used. Essay Submission 2/4=1/2 1/2*1/3=1/6 This essay has not been graded. Question 2 (Worth 10 points) (07.02 MC) Leena is arranging 3 different books in a row on a shelf. Create a sample space for the arrangement of a detective story (D), a mystery story (M), and a comic book (C). Essay Submission D,M,C M,D,C C,D,M C,M,D D,C,M M,C,D This essay has not been graded. Question 3 (Worth 10 points) (07.01 MC) A bag has 1 red marble, 4 blue marbles, and 3 green marbles. Peter draws a marble randomly from the bag, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer. Essay Submission He replaces it.So the probability of the first is going to be 4 out of 8 or 4/8 the second is the same because he replaced it so 4/8*4/8=16/64=8/32=4/16=2/8=1/4. This essay has not been graded. Question 4 (Worth 10 points) (07.03 MC) Chang has 2 shirts: a white one and a black one. He also has 2 pairs of pants, one blue and one tan. What is the probability, if Chang gets dressed in the dark, that he winds up wearing the white shirt and tan pants? Show your work. Essay Submission 1/2*1/2=1/4 This essay has not been graded. that

OpenStudy (anonymous):

Bloody 'ell

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

So what are you gonna do? Changed photo again?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Can you retake the exam?